# Multi-wavelength Emission from Jets and Magnetically Arrested Disks in Nearby Radio Galaxies: Application to M87

### v1

@@ Abstract
Radio galaxies are a subclass of active galactic nuclei that drive relativistic jets from their center and are observed in radio to very-high-energy gamma rays. The emission mechanisms and regions are still unknown. High-energy gamma rays can be explained by the emission from the magnetically arrested disks (MADs) around the central supermassive black hole, for which the magnetic flux threading the black hole is in a saturation level, although the emission from the MADs does not explain the optical and X-ray data. We construct a multi-wavelength emission model in which the optical and X-ray emission is emitted by jets and the gamma rays by MADs. Our model takes into account the particle injection by the magnetic reconnection at the jet base close to the black hole and particle entrainment from the ambient gas at the jet emission zone. We apply our model to M87 and find that our model can explain the simultaneous multi-wavelength data. In our model, the emission from the jets is the synchrotron radiation of the nonthermal electrons accelerated by magnetic reconnection, and the emission from the MADs is the synchrotron radiation mainly of the nonthermal protons accelerated by turbulence. We also find that the strong plasma entrainment is necessary to explain the multi-wavelength data. Our model will be tested by variability analysis among the multi-wavelength data.
@@ 1 Introduction
Leptonic jet models, in which energetic electrons radiate the multi-wavelength photons via synchrotron and inverse Compton scattering, are considered as the standard scenario (e.g., Abdo et al., 2009; MAGIC Collaboration et al., 2020). However, at least for M87, if we try to reproduce the observational data by the leptonic jet model, the magnetic field strength required to explain the multi-wavelength data is lower than that estimated by the observed core-shift in radio band (Kino et al., 2015; Jiang et al., 2021). If we use the core-shift-based magnetic field, the calculated gamma-ray flux is far below the observed gamma-ray flux (e.g., Lucchini et al., 2019; EHT MWL Science Working Group et al., 2021).
Kimura & Toma (2020) proposed a multi-wavelength emission model from the magnetically arrested disks (MADs; Bisnovatyi-Kogan & Ruzmaikin, 1974; Narayan et al., 2012) around the central supermassive black hole (BH), for which the magnetic flux threading the BH is in a saturation level. The MAD model can reproduce the gamma-ray data, but cannot explain the optical and X-ray data (Kimura & Toma, 2020; Kuze et al., 2022). This motivates us to construct a two-zone emission model in which the gamma-rays come from MADs and the optical and X-rays come from jets.
@@
In this paper, we construct a multi-wavelength emission model that takes into account both jets and MADs, which we call ‘Jet-MAD model’. We consider the plasma loading process into the jets in the MAD phase that makes the magnetization parameter of the jets depend on the BH mass and accretion rate (see Equations (6) and (7) below). In the jet, the magnetization parameter is much higher than unity, and we consider the dissipation mechanism at the dissipation region as magnetic reconnection. We also take into account the plasma entrainment from the ambient gas in that region.
@@ 2.1 MAD model
In this subsection, we briefly explain the particle acceleration and emission mechanisms of the MADs. We use the MAD model constructed by Kimura & Toma (2020) and Kuze et al. (2022) with modifications of heating and acceleration mechanisms. [^1] Kimura & Toma (2020) and Kuze et al. (2022) consider the magnetic reconnection as the heating and acceleration processes, while we consider the MHD turbulence dissipation. MHD turbulence in the accretion flows accelerates the particles (Lynn et al., 2014; Kimura et al., 2016, 2019; Sun & Bai, 2021), which emit multi-wavelength photons via synchrotron radiation.
We consider that the high-temperature plasma accretes onto the BH of mass $M_{\rm BH}$. The mass accretion rate, $\dot{M}$, and the size of the emission region, $R_{d}$, are normalized by the Eddington rate and the gravitational radius, respectively, i.e., $\dot{M}c^{2}=\dot{m}L_{\rm Edd}/c^{2}$, $R_{d}=rR_{g}=rGM_{\rm BH}/c^{2}$, where $L_{\rm Edd}$ is the Eddington luminosity, $c$ is the speed of light, and $G$ is the gravitational constant. We use the notation $Q_{X}=Q/10^{X}$ unless otherwise noted. For the BH mass, we use the notation $M_{9}=M_{\rm BH}/(10^{9}M_{\odot})$.
@@
Thermal electrons emit multi-wavelength photons via synchrotron and Comptonization processes. We determine the temperature of the thermal electrons by equating the heating rate with the cooling rate. We consider that the MHD turbulence heats up the thermal electrons. Kawazura et al. (2020) shows that when the compressive wave is dominant, the ratio of the electron-proton heating rate is approximated to be compressive-to-Alfvénic wave power ratio:
$\frac{Q_{i}}{Q_{e}}\approx\frac{P_{\rm compr}}{P_{\rm AW}},$ (1)
where $Q_{i}$ is the proton heating rate, $Q_{e}$ is the electron heating rate, $P_{\rm compr}$ is the compressive wave power, and $P_{\rm AW}$ is the Alfvénic wave power. The wave power ratio in the MAD has not been established yet. We give $Q_{i}/Q_{e}$ as a parameter, which is set to be 10. Since the accretion flow is in the MAD state, the plasma beta is likely lower than 1. Then, $Q_{i}/Q_{e}\simeq 10$ corresponds to $R_{\rm low}\simeq 10$ in the $R_{\rm low}-R_{\rm high}$ prescription (see Mościbrodzka et al., 2016), and $R_{\rm low}\simeq 10$ is favored by the polarization observation of M87 (Event Horizon Telescope Collaboration et al., 2021b). The cooling rates of the thermal electrons are calculated in the same manner as Kimura et al. (2015). For M87, the accretion rate is low, and then, the bremsstrahlung is inefficient as a cooling process because of the low electron number density.
Primary electrons and protons are accelerated by the MHD turbulence and emit nonthermal synchrotron radiation. The energy distribution of nonthermal particles is required to calculate the photon spectra. To obtain the nonthermal energy distribution, we solve the energy transport equation by assuming the one-zone and steady-state approximations (see Equation (3) of Kimura & Toma, 2020). We use the injection term as a power-law form whose index is $s_{\rm inj}$ with an exponential cutoff as in Equation (5) of Kimura & Toma (2020). The maximum energy of the particles is determined by equating the acceleration timescale with the energy loss timescale.
$D_{pp}\approx\frac{p^{2}}{3\eta_{\rm turb}}\left(\frac{c}{H}\right)\left(\frac{V_{A}}{c}\right)^{2},$ (2)
where $V_{A}\approx B_{d}/\sqrt{4\pi\rho}$ is the Alfvén velocity, $B_{d}\approx\sqrt{8\pi\rho C_{s}^{2}/\beta}$ is the magnetic field in the MADs, $\rho\approx\dot{M}/(4\pi R_{d}HV_{R})$ is the mass density, $H\sim(C_{s}/V_{K})R_{d}$ is the scale height of the MADs, $C_{s}\approx V_{K}/2$ is the sound speed, $V_{R}\approx\alpha V_{K}$ is the radial velocity, $V_{K}=\sqrt{GM_{\rm BH}/R_{d}}$ is the Kepler velocity, $\alpha=0.3$ is the viscous parameter, $\beta=0.1$ is the plasma beta, $\eta_{\rm turb}H$ is the effective mean free path, and $\eta_{\rm turb}=0.833$ is the numerical factor (see also Kimura & Toma, 2020; Kuze et al., 2022). Then, we can write the acceleration timescale as
$t_{\rm acc}=\frac{p^{2}}{D_{pp}}\approx 3\eta_{\rm turb}\frac{H}{c}\left(\frac{c}{V_{A}}\right)^{2}.$ (3)
We take into account relevant cooling processes (synchrotron, Bethe-Heitler, photomeson production), diffusive escape, and infall to the BH as the loss processes. For diffusive escape, we estimate the timescale as $t_{\rm diff}=R_{d}^{2}/D_{zz}$, where $D_{zz}=(1/3)\eta_{\rm turb}cH$ is the diffusion coefficient in space.
@@ 2.2 Physical quantities at the base of the jet
The magnetic reconnection in the BH magnetosphere is expected to produce the power-law electron energy distribution with the index $p_{\rm inj}\sim 1$ (e.g., Sironi & Spitkovsky, 2014; Guo et al., 2014, 2015, 2016; Werner et al., 2016). These injected electrons rapidly cool as the cooling timescale is much shorter than the dynamical timescale (Kimura et al., 2022), resulting in the index of nonthermal electrons, $p\sim 2$. Then, the synchrotron spectrum is power-law with the index $(p-1)/2\approx 0.5$ up to $E_{\gamma,\rm cut,ms}=\gamma_{\rm cut,ms}^{2}heB_{\rm rec}/(2\pi m_{e}c)$, where $\gamma_{\rm cut,ms}$ is the cutoff Lorentz factor achieved by the magnetic reconnection in the BH magnetosphere, $B_{\rm rec}$ is the magnetic field in the reconnection region, $h$ is the Planck constant, $e$ is the electron charge, and $m_{e}$ is the electron mass. We assume the cutoff Lorentz factor as
$\gamma_{\rm cut,ms}=4\sigma_{\pm},$ (4)
where $\sigma_{\pm}$ is the magnetization parameter. The synchrotron photon spectrum has an exponential cutoff for photons higher than $E_{\gamma,\rm cut,ms}$.
$n_{\pm}=\dot{n}_{\pm}\frac{l_{\rm rec}}{c},$ (5)
@@
where $\dot{n}_{\pm}$ is the pair production rate, which depends on $\sigma_{\pm}$ through Equation (4), and $l_{\rm rec}$ is the size of the reconnection region. According to Chen et al. (2023), we estimate the steady-state magnetization parameter by solving
$\sigma_{\pm}=\frac{B_{\rm rec}^{2}}{4\pi n_{\pm}m_{e}c^{2}}.$ (6)
@@
The steady-state magnetization parameter depends only on $B_{\rm rec}$ and $l_{\rm rec}$. We consider that the accretion flow is in the MAD state, and magnetic field strength around the BH is estimated to be $B_{H}=\sqrt{\dot{M}c\Phi^{2}_{H}/(4\pi^{2}R_{g}^{2})}\simeq 1.1\times 10^{3}M_{9}^{-1/2}\dot{m}_{-4}^{1/2}\Phi_{H,1.7}\ \rm G$, where $\Phi_{H}\simeq 50\Phi_{H,1.7}$ is the saturated magnetic flux of the MAD. High-resolution GRMHD simulation with a BH spin parameter $a=0.935$ suggests that the magnetic reconnection occurs in a radius $r_{\rm rec}\approx 2R_{g}$ (Ripperda et al., 2022), and we fix $r_{\rm rec}=2R_{g}$. We estimate $B_{\rm rec}$ in the same way as Kimura et al. (2022) and to be $B_{\rm rec}=\sqrt{2}B_{H}\left(r_{\rm rec}/{R_{g}}\right)^{-2}\simeq 3.9\times 10^{2}M_{9}^{-1/2}\dot{m}_{-4}^{1/2}\Phi_{\rm rec,1.2}\ \rm G$, where $\Phi_{\rm rec}=\sqrt{2}\Phi_{H}(r_{\rm rec}/R_{g})^{-2}$ is the magnetic flux in the magnetic reconnection region. We fix $l_{\rm rec}=R_{g}$ for simplicity.
Figure 2: Magnetization parameters for BH mass of $10^{9}M_{\odot}$ as a function of $\dot{m}$, calculated in the methods of Chen et al. (2023) (red solid line) and Kimura et al. (2022) (blue dashed line) and estimated by Equation (7) (green dash-dotted line).
We show $\sigma_{\pm}$ for the BH mass of $10^{9}M_{\odot}$ as a function of $\dot{m}$ in Figure 2. As shown in Figure 2, $\sigma_{\pm}$ approximately depends on $\dot{m}^{-1/4}$. This comes from the self-regulation of the specific luminosity to peak at $E_{\gamma,\rm cut,ms}\approx 0.5m_{e}c^{2}$. Then, we analytically estimate $\sigma_{\pm}$ as
$\sigma_{\pm}\simeq 5.9\times 10^{4}M_{9}^{1/4}\dot{m}_{-4}^{-1/4}\Phi_{\rm rec,1.2}^{-1/2}.$ (7)
Kimura et al. (2022) also estimate the magnetization parameter after pair loading (see their Equation (11)), but our estimate of $\sigma_{\pm}$ differs from that given in Kimura et al. (2022). They assume photon spectrum with an abrupt cutoff (i.e., no photons above $E_{\gamma,\rm cut,ms}$), which leads to different values of $\sigma_{\pm}$ and conclusions for such a range of mass accretion rates that we consider in this paper. [^2] There is another point that differs between ours and that given in Kimura et al. (2022). Ours takes into account the feedback by the pair loading when $m_{e}c^{2}<E_{\gamma,\rm cut,ms}<E_{\gamma,\rm rad}$, where $E_{\gamma,\rm rad}$ is the synchrotron photon energy for the burn-off limit (see Equation (5) in Kimura et al. (2022)), which help achieving self-regulation given in Equation (7). In contrast, for the cases with a mass accretion rate as low as that in Sgr A*, the estimate given in Kimura et al. (2022) should be valid, as neither the feedback nor the tail distribution affects the pair loading process. Our model of pair loading assumes that $E_{\gamma,\rm cut,ms}$ determines the cutoff energy because of the sufficient pair loading, which contradicts the situation of low accretion rate cases where $E_{\gamma,\rm rad}<E_{\gamma,\rm cut,ms}$ is satisfied.
@@ 2.3 Energy contents of the jet
In the magnetically dominated jet region, the magnetic field is dominated by the toroidal component and $B\propto R^{-1}$ at $R\gtrsim 10R_{g}$, where $R$ is the distance from the jet axis (Kimura et al., 2022). If all the electron-positron pairs inside the blob are accelerated up to $\sim c$ by efficient electromagnetic bulk acceleration, number flux conservation results in $n_{e}\propto R^{-2}$. Then, the magnetization parameter is conserved inside the jet, i.e., $\sigma\propto B^{2}/n_{e}\propto R^{0}$. However, the bulk acceleration is not efficient as observed in the radio jet of M87 (see Figure 12 of Park et al., 2019). We consider that the blob expands due to the velocity dispersion of electron-positron pairs, along the magnetic field lines, which decreases the number density more rapidly than $n_{e}\propto R^{-2}$. By employing a numerical factor $\xi$, we write the magnetization parameter at the dissipation region before dissipation as $\xi\sigma_{\pm}$.
The interval time of the magnetic reconnection flares in the BH magnetosphere is around $\Delta t\simeq 10^{3}R_{g}/c$ (Ripperda et al., 2022), which is comparable to the viscous timescale of the MADs. Thus, we consider that the blobs will expand $\sim 10^{3}$ times along the magnetic field lines. In this paper, we set $\xi=10^{3}$.
@@
$L_{j}$ $=\Gamma^{2}\left(n_{e}m_{e}c^{2}+\frac{B^{2}}{8\pi}\right)\pi R_{j}^{2}\beta c$
$=\Gamma^{2}(1+0.5\xi\sigma_{\pm})n_{e}m_{e}c^{2}\pi R_{j}^{2}\beta c,$ (9)
where $R_{j}$ is the size of the dissipation region, which we treat as a free parameter, $\Gamma$ is the jet Lorentz factor, and $\beta$ is the jet speed normalized by $c$. We estimate $R_{j}$ and $\Gamma$ later in this subsection.
At the dissipation region, the magnetization parameter is $\xi\sigma_{\pm}\gg 1$, so that converting the magnetic energy to the thermal energy is more efficient than converting the plasma kinetic energy. We consider that the velocity difference between the jet and the ambient gas triggers the Kelvin-Helmholtz instability (KHI) at the edge of the jet (Sironi et al., 2021; Davelaar et al., 2023), inducing the turbulence. As a result, the electron-proton plasma can be entrained into the jet from the ambient gas, which consists of the wind from the radiatively inefficient accretion flow (RIAF; Narayan & Yi, 1994; Yuan & Narayan, 2014) in the outer region of the MAD (see right part of Figure 1). The KHI-induced turbulence also causes magnetic reconnection inside the jet. Thus, the magnetization parameter at the dissipation region decreases due to the plasma entrainment and the dissipation of the magnetic energy through the magnetic reconnection (e.g., Hoshino & Lyubarsky, 2012; Guo et al., 2020; Sironi et al., 2021; Mehlhaff et al., 2024). Such turbulent reconnection accelerates nonthermal particles efficiently (e.g., Hoshino, 2012; Xu & Lazarian, 2023). We define the magnetization parameter after the plasma entrainment $\sigma_{\rm ent}$ as a free parameter. We also define the magnetization parameter after dissipation as $\sigma_{\rm dis}$.
@@
$L_{j}$ $=\Gamma^{2}\left(n_{e}m_{e}c^{2}+n_{p}m_{p}c^{2}+\frac{B^{2}}{8\pi}\right)\pi R_{j}^{2}\beta c$
$=\Gamma^{2}(1+0.5\sigma_{\rm ent})(n_{e}m_{e}c^{2}+n_{p}m_{p}c^{2})\pi R_{j}^{2}\beta c,$ (10)
where $n_{p}$ is the proton number density. If the plasma entrainment is strong ($n_{e}\approx n_{p}$), the sum of the electron and proton energy density is written as $n_{e}m_{e}c^{2}+n_{p}m_{p}c^{2}\sim n_{p}m_{p}c^{2}$.
@@
$L_{j}$ $=L_{e}+L_{p}+\Gamma^{2}\frac{B_{\rm dis}^{2}}{8\pi}\pi R_{j}^{2}\beta c$
@@
We estimate the jet Lorentz factor and Doppler factor as follows. The very long baseline interferometry (VLBI) observations for M87 estimate the jet width as $R\approx z^{2/3}$, where $z$ is the distance from the BH (Hada et al., 2013; Nakamura et al., 2018). Thus, we assume the size of the dissipation region as $R_{j}=10^{2}R_{g}$ to have the distance of the dissipation region $z_{j}\simeq 10^{3}R_{g}$. Hereafter, we use the notation $R_{j,2}=R_{j}/(10^{2}R_{g})$. We set the jet Lorentz factor $\Gamma$ at the dissipation region as $\Gamma\beta\simeq 0.8$ based on the observation of the superluminal motion of the M87 jet with KaVA and VLBA (Park et al., 2019). The M87 jet inclination angle is observed as $\theta_{j}\approx 17\ \rm deg$ by the observation of jet and counter jet proper motions (Walker et al., 2018). Hence, we estimate the Doppler factor to be $\delta_{D}\approx 1.9$. The variability timescale is then estimated to be $t_{v}=R_{j}/(c\delta_{D})\simeq 1.6\times 10^{6}\;R_{j,2}\;{\rm s}$, which is consistent with the observed X-ray daily variability (Imazawa et al., 2021; EHT MWL Science Working Group et al., 2021). We calculate the photon spectra observed on Earth by $\nu_{\rm obs}L_{\nu,\rm obs}=\delta_{D}^{4}\nu_{\rm src}L_{\nu,\rm src}$, where $\nu_{\rm src}L_{\nu,\rm src}$ is the intrinsic photon spectrum obtained by the jet model.

### v2

@@ Abstract
Radio galaxies are a subclass of active galactic nuclei that drive relativistic jets from their center and are observed in radio to very-high-energy gamma rays. The emission mechanisms and regions are still unknown. High-energy gamma rays can be explained by the emission from the magnetically arrested disks (MADs) around the central supermassive black hole, for which the magnetic flux threading the black hole is in a saturation level, although the emission from the MADs does not explain the optical and X-ray data. We construct a two-zone multi-wavelength emission model in which optical and X-rays come from jets, while mm/sub-mm and gamma rays come from MADs. Our model takes into account the particle injection by the magnetic reconnection at the jet base close to the black hole and particle entrainment from the ambient gas at the jet emission zone. We apply our model to M87 and find that our model can explain the simultaneous multi-wavelength data, except for the radio data which could be explained if we extend our one-zone emission model to a one-dimensional one. We also find that the strong plasma entrainment is necessary to explain the multi-wavelength data. Our model will be tested by variability analysis among the multi-wavelength data.
@@ 1 Introduction
Leptonic jet models, in which energetic electrons radiate the multi-wavelength photons via synchrotron and inverse Compton scattering, are considered as the standard scenario (e.g., Abdo et al., 2009; MAGIC Collaboration et al., 2020). However, at least for M87, if we try to reproduce the observational data by the leptonic jet model, the magnetic field strength required to explain the multi-wavelength data ($\sim$ 1-10 mG) is lower than that estimated by the observed core-shift in radio band ($\sim$ 1 G)(Kino et al., 2015; Jiang et al., 2021). If we use the core-shift-based magnetic field, the calculated gamma-ray flux is far below the observed gamma-ray flux (e.g., Lucchini et al., 2019; EHT MWL Science Working Group et al., 2021).
Kimura & Toma (2020) proposed a multi-wavelength emission model from the magnetically arrested disks (MADs; Bisnovatyi-Kogan & Ruzmaikin, 1974; Narayan et al., 2012) around the central supermassive black hole (BH), for which the magnetic flux threading the BH is in a saturation level. The MAD model can reproduce the mm/sub-mm and gamma-ray data, but cannot explain the optical and X-ray data (Kimura & Toma, 2020; Kuze et al., 2022). This motivates us to construct a two-zone emission model in which the mm/sub-mm and gamma-rays come from MADs and the optical and X-rays come from jets.
@@
In this paper, we construct a two-zone multi-wavelength emission model that takes into account both jets and MADs, which we call ‘Jet-MAD model’. We consider the plasma loading process into the jets in the MAD phase that makes the magnetization parameter of the jets depend on the BH mass and accretion rate (see Equation (1) and Appendix B). In the jet, the magnetization parameter is much higher than unity, and we consider the dissipation mechanism at the dissipation region as magnetic reconnection. We also take into account the plasma entrainment from the ambient gas in that region.
@@ 2.1 MAD model
In this subsection, we briefly explain the particle acceleration and emission mechanisms of the MADs. We use the one-zone emission model, called ‘MAD model’, constructed by Kimura & Toma (2020) and Kuze et al. (2022) with modifications of heating and acceleration mechanisms. [^1] Kimura & Toma (2020) and Kuze et al. (2022) consider the magnetic reconnection as the heating and acceleration processes, while we consider the MHD turbulence dissipation. We explain the detailed calculation method of the MAD model in Appendix A. We consider that the high-temperature plasma accretes onto the BH of mass $M_{\rm BH}$. The mass accretion rate, $\dot{M}$, and the size of the emission region, $R_{d}$, are normalized by the Eddington rate and the gravitational radius, respectively, i.e., $\dot{M}c^{2}=\dot{m}L_{\rm Edd}$, $R_{d}=rR_{g}=rGM_{\rm BH}/c^{2}$, where $L_{\rm Edd}$ is the Eddington luminosity, $c$ is the speed of light, and $G$ is the gravitational constant. We use the notation $Q_{X}=Q/10^{X}$ unless otherwise noted. For the BH mass, we use the notation $M_{9}=M_{\rm BH}/(10^{9}M_{\odot})$.
@@
Thermal electrons emit multi-wavelength photons via the synchrotron and Comptonization processes. We determine the thermal electron temperature by equating the electron heating rate with the cooling rate. We consider that the MHD turbulence heats up the thermal electron. Electron heating processes by collisionless MHD turbulence in the MADs are still under debate (cf. Kawazura et al., 2020). Then, we give the ratio of electron-proton heating rate, $Q_{i}/Q_{e}$, as a parameter, and we set $Q_{i}/Q_{e}\simeq 10$. The cooling rates of the thermal electrons are calculated in the same manner as Kimura et al. (2015). For M87, the accretion rate is low, and then, the bremsstrahlung is inefficient as a cooling process because of the low electron number density.
MHD turbulence accelerates the primary electrons and protons which emit nonthermal synchrotron radiation (Lynn et al., 2014; Kimura et al., 2016, 2019b; Sun & Bai, 2021). To obtain the energy distributions of the nonthermal particles, we solve the energy transport equation with the one-zone and steady-state approximations, and we use the injection term as a power-law form with an exponential cutoff (see Appendix A). We estimate the cutoff energy of the particles by equating the acceleration timescale with the loss timescale.
As the acceleration process, we consider stochastic acceleration by the MHD turbulence with a hard-sphere-like diffusion coefficient. Within the diffusion process, we write the effective mean free path as $\eta_{\rm turb}H$, where $H$ is the scale height of the accretion flow and $\eta_{\rm turb}$ is the numerical factor. We consider the relevant cooling processes (synchrotron, Bethe-Heitler, photomeson production), diffusive escape, and infall to the BH for the loss processes.
@@ 2.2 Physical quantities at the base of the jet
In this paper, we estimate the steady-state magnetization parameter in the reconnection region by equating the pair production rate with the pair escape rate (see Chen et al., 2023). We show the detailed calculation model of the steady-state magnetization parameter in Appendix B. According to Chen et al. (2023), we estimate the steady-state magnetization parameter by solving
$\sigma_{\pm}=\frac{B_{\rm rec}^{2}}{4\pi n_{\pm}m_{e}c^{2}},$ (1)
where $\sigma_{\pm}$ is the magnetization parameter, $B_{\rm rec}$ is the magnetic field in the reconnection region, $n_{\pm}$ is the produced pair number density in the reconnection region, and $m_{e}$ is the electron mass.
@@
As $n_{\pm}$ depends on $\sigma_{\pm}$ and the size of the reconnection region, $l_{\rm rec}$, the steady-state magnetization parameter depends only on $B_{\rm rec}$ and $l_{\rm rec}$ (see Appendix B). We consider that the accretion flow is in the MAD state, and magnetic field strength around the BH is estimated to be $B_{H}=\sqrt{\dot{M}c\Phi^{2}_{H}/(4\pi^{2}R_{g}^{2})}\simeq 1.1\times 10^{3}M_{9}^{-1/2}\dot{m}_{-4}^{1/2}\Phi_{H,1.7}\ \rm G$, where $\Phi_{H}\simeq 50\Phi_{H,1.7}$ is the saturated magnetic flux of the MAD. High-resolution GRMHD simulation with a BH spin parameter $a=0.935$ suggests that the magnetic reconnection occurs in a radius $r_{\rm rec}\approx 2R_{g}$ (Ripperda et al., 2022), and we fix $r_{\rm rec}=2R_{g}$. We estimate $B_{\rm rec}$ in the same way as Kimura et al. (2022) and to be $B_{\rm rec}=\sqrt{2}B_{H}\left({r_{\rm rec}}/{{R_{g}}}\right)^{-2}\simeq 3.9\times 10^{2}M_{9}^{-1/2}\dot{m}_{-4}^{1/2}\Phi_{\rm rec,1.2}\ \rm G$, where $\Phi_{\rm rec}=\sqrt{2}\Phi_{H}(r_{\rm rec}/R_{g})^{-2}$ is the magnetic flux in the magnetic reconnection region. We fix $l_{\rm rec}=R_{g}$ for simplicity.
@@ 2.3 Energy contents of the jet
In the magnetically dominated jet region, the magnetic field is dominated by the toroidal component and $B\propto R^{-1}$ at $R\gtrsim 10R_{g}$, where $R$ is the distance from the jet axis (Kimura et al., 2022). If all the electron-positron pairs inside the blob are accelerated up to $\sim c$ by efficient electromagnetic bulk acceleration, number flux conservation results in $n_{e}\propto R^{-2}$. Then, the magnetization parameter is conserved inside the jet, i.e., $\sigma\propto B^{2}/n_{e}\propto R^{0}$. However, the bulk acceleration is not efficient as observed in the radio jet of M87 (see Figure 12 of Park et al., 2019). We consider that the single, one-zone blob expands due to the velocity dispersion of electron-positron pairs, along the magnetic field lines, which decreases the number density more rapidly than $n_{e}\propto R^{-2}$. By employing a numerical factor $\xi$, we write the magnetization parameter at the dissipation region before dissipation as $\xi\sigma_{\pm}$.
The interval time of the magnetic reconnection flares in the BH magnetosphere is around $\Delta t\simeq 10^{3}R_{g}/c$ (Ripperda et al., 2022), which is comparable to the viscous timescale of the MADs. Thus, we consider that the blob will expand $\sim 10^{3}$ times along the magnetic field lines. In this paper, we set $\xi=10^{3}$.
@@
$L_{j}$ $=\Gamma_{j}^{2}\left(n_{e}m_{e}c^{2}+\frac{B^{2}}{8\pi}\right)\pi R_{j}^{2}\beta_{j}c$
$=\Gamma_{j}^{2}(1+0.5\xi\sigma_{\pm})n_{e}m_{e}c^{2}\pi R_{j}^{2}\beta_{j}c,$ (3)
where $R_{j}$ is the size of the dissipation region, which we treat as a free parameter, $\Gamma_{j}$ is the jet Lorentz factor, and $\beta_{j}$ is the jet speed normalized by $c$. We estimate $R_{j}$ and $\Gamma_{j}$ in Appendix C. Hereafter, we use the notation $R_{j,2}=R_{j}/(10^{2}R_{g})$.
At the dissipation region, the magnetization parameter is $\xi\sigma_{\pm}\gg 1$, so that converting the magnetic energy to the thermal energy is more efficient than converting the plasma kinetic energy. We consider that the velocity difference between the jet and the ambient gas triggers the Kelvin-Helmholtz instability (KHI) at the edge of the jet (Sironi et al., 2021; Davelaar et al., 2023), inducing the turbulence. As a result, the electron-proton plasma can be entrained into the jet from the ambient gas, which consists of the wind from the radiatively inefficient accretion flow (RIAF; Narayan & Yi, 1994; Yuan & Narayan, 2014) in the outer region of the MAD (see right part of Figure 1). The KHI-induced turbulence also causes magnetic reconnection inside the jet. Thus, the magnetization parameter at the dissipation region decreases due to the plasma entrainment and the dissipation of the magnetic energy through the magnetic reconnection (e.g., Hoshino & Lyubarsky, 2012; Guo et al., 2020; Sironi et al., 2021; Mehlhaff et al., 2024). Such turbulent reconnection accelerates nonthermal particles efficiently (e.g., Hoshino, 2012; Xu & Lazarian, 2023). We define the magnetization parameter after the plasma entrainment $\sigma_{\rm ent}$ as a free parameter. The parameter, $\sigma_{\rm ent}$, determines the amount of the plasma entrainment. If the plasma entrainment is strong(weak), it results in low(high) $\sigma_{\rm ent}$. We also define the magnetization parameter after dissipation as $\sigma_{\rm dis}$.
@@
$L_{j}$ $=\Gamma_{j}^{2}\left(n_{e}m_{e}c^{2}+n_{p}m_{p}c^{2}+\frac{B^{2}}{8\pi}\right)\pi R_{j}^{2}\beta_{j}c$
$=\Gamma_{j}^{2}(1+0.5\sigma_{\rm ent})(n_{e}m_{e}c^{2}+n_{p}m_{p}c^{2})\pi R_{j}^{2}\beta_{j}c,$ (4)
where $n_{p}$ is the proton number density and $m_{p}$ is the proton mass. If the plasma entrainment is strong ($n_{e}\approx n_{p}$), the sum of the electron and proton energy density is written as $n_{e}m_{e}c^{2}+n_{p}m_{p}c^{2}\sim n_{p}m_{p}c^{2}$.
@@
$L_{j}$ $=L_{e}+L_{p}+\Gamma_{j}^{2}\frac{B_{\rm dis}^{2}}{8\pi}\pi R_{j}^{2}\beta_{j}c$